40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 732     Entry time: Thu Jul 24 03:08:20 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: +f2 DRMI+2ARMS 

rob, john, yoichi

Tonight we tried to move the 166MHz (f2) sideband frequency by changing the settings on the Marconi. Reducing the frequency by 4kHz reduced the amplitude of the 166MHz sidebands, but we were still able to lock the DRMI with the +-f2 sidebands by electronically compensating for the gain decrease, and also to lock the DRMI+2ARMs while resonating the -f2 sideband. No luck with the +f2.

Then we larkily tried increasing the frequency by 4kHz, which ~doubled the f2 sideband transmission through the MC. This means our frequencies/MC length have been mismatched for months. Apparently I explained the level of the f2 sidebands by just imagining that I'd (or someone) had set the modulation depth at that level some time in the past.

It's a miracle any locking worked at all in this state. Once this was done and we worked out a few kinks in the script, adjusting some gains to compensate, we managed to get the DRMI+2ARMS to lock a couple of times while resonating the +f2 sideband. It takes a while, but at least it happens. Tomorrow we'll measure the length of the mode cleaner properly and then try again. No need to vent just yet.
ELOG V3.1.3-